# Knockoffs for low dimensions: changing the nominal level post-hoc to gain power while controlling the FDR

### v1

@@ 3.1 Increase $\alpha$ to gain power
where the parameter $\alpha^{\textnormal{kn}}\in(0,1]$ can be interpreted as the prespecified initial significance level. The difference of $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}$ to the stopping time $T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}$ of Barber and Candès [1] (see equation (7)) is the second condition $\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-t\}=0$. Be aware that this condition only comes into play if the first condition is not satisfied for any $t$, meaning $T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}=T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}$ if $R^{\textnormal{BC}}\neq\emptyset$. However, in case of $T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}=\infty$, we have $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}<\infty$. Similar to BC-knockoff, our procedure, which we call post-hoc knockoff (ph-knockoff), rejects all hypotheses $H_{0}^{i}$ with $W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}$. Since $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\leq T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}$, ph-knockoff is uniformly more powerful than BC-knockoff by Barber and Candès [1]. However, in case of $T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}=\infty$, we can only report the rejections at a larger significance level $\alpha>\alpha^{\textnormal{kn}}$, as we show in the next section (see Theorem 3.1).
Remark 1.
Ren and Barber [21] introduced a modification of the stopping time $T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}$ that is similar to $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}$, adding the condition $\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq t\}<1/\alpha^{\textnormal{kn}}$ instead of $\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-t\}=0$ (see Remark 5 in [21]). They used this adjusted stopping time to improve their derandomized knockoff procedure (see also Section 4) without considering a post-hoc choice of $\alpha$. While our approach of switching $\alpha$ post-hoc would also work with their stopping time, we think our stopping time leads to smaller post-hoc levels.
@@ 3.2 Decrease $\alpha$ to improve precision
In Algorithm 10, we summarize our general procedure, where we denote by $\alpha^{\textnormal{ph}}$ our post-hoc significance level and by $R^{\textnormal{ph}}$ the corresponding rejection set. If BC-knockoff does not make any rejections, we increase the significance level ($\alpha^{\textnormal{ph}}>\alpha^{\textnormal{kn}}$), and thus potentially obtain rejections. If BC-knockoff makes rejections, we decrease the significance level ($\alpha^{\textnormal{ph}}\leq\alpha^{\textnormal{kn}}$) to increase the evidence. In this way, the claims made by ph-knockoff are always stronger than the claims made by BC-knockoff.
@@
Output: Rejection set $R^{\textnormal{ph}}$ and post-hoc level $\alpha^{\textnormal{ph}}$.
@@
4:  $\alpha^{\textnormal{ph}}=\frac{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}}{\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}}$
@@
6:  $\alpha^{\textnormal{ph}}=\alpha^{\textnormal{kn}}$
@@
8:return $R^{\textnormal{ph}}$, $\alpha^{\textnormal{ph}}$
@@
The validity of these local e-values (i.e. that $\mathbb{E}_{H_{0}^{S}}[E_{S}^{\textnormal{ph}}]\leq 1$) follows since $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}$ is a stopping time with respect to the filtration generated by a masked version of $(W_{i})_{i\in[p]}$, meaning the event $\{T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\leq t\}$ is determined by the absolute values $|W_{i}|$ for all $i$ and $\sum_{i:|W_{i}|\geq s}\mathbbm{1}\{W_{i}>0\}$ for all $s\leq t$ (see Remark 4 in [21]).
Theorem 3.1.
@@
$\alpha^{\textnormal{ph}}=\frac{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}}{\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}}\quad\text{with the convention }\alpha^{\textnormal{ph}}=\alpha^{\textnormal{kn}}\text{ if }\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}=0.$ (13)
Then $R^{\textnormal{ph}}\in\mathcal{R}_{\alpha^{\textnormal{ph}}}(\boldsymbol{E}^{\textnormal{ph}})$, where $R^{\textnormal{ph}}=\{i\in[p]:W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}$. Therefore, it holds that
$\mathbb{E}\left[\frac{\textnormal{FDP}_{I_{0}}(R^{\textnormal{ph}})}{\alpha^{\textnormal{ph}}}\right]\leq 1.$ (14)
Furthermore, $R^{\textnormal{ph}}\supseteq R^{\textnormal{BC}}$, meaning ph-knockoff always makes as least as many rejections as BC-knockoff and $\alpha^{\textnormal{ph}}\leq\alpha^{\textnormal{kn}}$ on the event $\{R^{\textnormal{BC}}\neq\emptyset\}$, meaning if BC-knockoff rejects any hypothesis then ph-knockoff does not report a larger significance level.
Proof.
We need to show that $E_{S}^{\textnormal{ph}}\geq\textnormal{FDP}_{S}(R^{\textnormal{ph}})/\alpha^{\textnormal{ph}}$ for all $S\subseteq[p]$. First, note that in case of $\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}=0$, we have $R^{\textnormal{ph}}=\emptyset$ and therefore the condition is trivially fulfilled. Hence, assume $R^{\textnormal{ph}}\neq\emptyset$. Then,
$E_{S}^{\textnormal{ph}}=\frac{|S\cap R^{\textnormal{ph}}|}{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}}=\frac{\textnormal{FDP}_{S}(R^{\textnormal{ph}})|R^{\textnormal{ph}}|}{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}}=\frac{\textnormal{FDP}_{S}(R^{\textnormal{ph}})}{\alpha^{\textnormal{ph}}}.$ (15)
Remark 2.
Note that $\alpha^{\textnormal{ph}}$ equals 11 in case of $\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}=1$. While rejections at a level of 11 seems like a trivial claim in the fixed-$\alpha$ regime, it is not for random $\alpha$ [14]. However, if such large values for $\alpha^{\textnormal{ph}}$ should be avoided, one could just set $\alpha^{\textnormal{ph}}=\alpha^{\textnormal{kn}}$ and $R^{\textnormal{ph}}=\emptyset$ in those cases. This obviously maintains (14).
@@ 4 Post-hoc $\alpha$ with derandomized knockoffs
Ren and Barber [21] introduced a derandomized knockoff procedure, which we call RB-knockoff in the following, by generating multiple knockoff copies and merging the results using an average of (compound) e-values. In the following, we briefly recap their method and demonstrate how it can be improved by using a post-hoc significance level. It should be noted that we present the RB-knockoff procedure with our stopping time $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}$ to be consistent with the previous part, even though Ren and Barber [21] proposed a slightly different one (see Remark 1). However, our approach generally works for any stopping time.
@@
Ren and Barber [21] showed that $E_{1}^{\text{avg}},\ldots,E_{p}^{\text{avg}}$ are compound e-values, meaning $\mathbb{E}\left[\sum_{i\in I_{0}}E_{i}^{\text{avg}}\right]\leq p$. The RB-knockoff procedure is defined by e-BH [32] applied on $E_{1}^{\text{avg}},\ldots,E_{p}^{\text{avg}}$ at prespecified level $\alpha^{\textnormal{ebh}}\in(0,1]$ and controls the FDR at level $\alpha^{\textnormal{ebh}}$.
In case of $k=1$ and $\alpha^{\textnormal{ebh}}=\alpha^{\textnormal{kn}}$, RB-knockoff coincides with BC-knockoff [21]. However, Ren and Barber [21] showed that $\alpha^{\textnormal{ebh}}=\alpha^{\textnormal{kn}}$ often leads to zero power in case of $k>1$ and therefore recommend to choose $\alpha^{\textnormal{ebh}}>\alpha^{\textnormal{kn}}$. The choice of the parameter $\alpha^{\textnormal{ebh}}$ is crucial. If $\alpha^{\textnormal{ebh}}$ is too small, one might not be able to make any rejections even though the signal is strong. If $\alpha^{\textnormal{ebh}}$ is too large, the results become imprecise. Ren and Barber [21] derive some concrete choices like $\alpha^{\textnormal{ebh}}=2\alpha^{\textnormal{kn}}$ based on simulations and heuristics, but they do not propose an optimal parameter. We note that it is possible to choose $\alpha^{\textnormal{ebh}}$ completely data-dependent while maintaining (10). In this way, we can just choose the $\alpha^{\textnormal{ebh}}$ that leads to the most desirable rejection set.
Theorem 4.1.
@@
$\mathbb{E}\left[\sup_{\alpha^{\textnormal{ebh}}\in(0,1]}\frac{\textnormal{FDP}_{I_{0}}(R_{\alpha^{\textnormal{ebh}}}^{\textnormal{RB}})}{\alpha^{\textnormal{ebh}}}\right]\leq 1.$ (17)
@@
The proof of the Theorem 4.1 follows immediately from the equivalence of e-BH with compound e-values $E_{1}^{\text{avg}},\ldots,E_{p}^{\text{avg}}$ and the e-Closure Principle with the following local e-values (see Theorem 23 in [36])
@@
There are multiple reasonable choices to pick $\alpha^{\textnormal{ebh}}$ based on the data that could lead to different rejection sets (in the following we refer to our data-dependent choice as $\alpha^{\textnormal{dph}}$). In order to compare our derandomized ph-knockoff procedure to RB-knockoff in the following, we start with some initial level $\alpha^{\textnormal{ebh}}\in(0,1)$ that should be chosen as in the RB-knockoff procedure. If RB-knockoff does not make any rejections with that $\alpha^{\textnormal{ebh}}$, we choose $\alpha^{\textnormal{dph}}>\alpha^{\textnormal{ebh}}$ as small as possible such that $R_{\alpha^{\textnormal{dph}}}^{\textnormal{RB}}$ is not empty. If $R_{\alpha^{\textnormal{ebh}}}^{\textnormal{RB}}\neq\emptyset$, then we choose $\alpha^{\textnormal{dph}}\leq\alpha^{\textnormal{ebh}}$ as small as possible while $R_{\alpha^{\textnormal{dph}}}^{\textnormal{RB}}=R_{\alpha^{\textnormal{ebh}}}^{\textnormal{RB}}$. In this way, our derandomized ph-knockoff procedure always rejects as least as many hypotheses as RB-knockoff. Furthermore, if RB-knockoff makes rejections, then $\alpha^{\textnormal{dph}}\leq\alpha^{\textnormal{ebh}}$. Our general procedure is summarized in Algorithm 2.
@@
Output: Rejection set $R^{\textnormal{dph}}$ and post-hoc level $\alpha^{\textnormal{dph}}$.
@@
5:  $\alpha^{\textnormal{dph}}=p/(i^{\textnormal{ebh}}E_{(i^{\textnormal{ebh}})}^{\text{avg}})$
@@
10:   $\alpha^{\textnormal{dph}}=\alpha^{\textnormal{kn}}$
@@
13:   $\alpha^{\textnormal{dph}}=p/(i^{*}E_{(i^{*})}^{\text{avg}})$
14:   $R^{\textnormal{dph}}=\{i\in[p]:E_{i}^{\text{avg}}\geq p/(i^{*}\alpha^{\textnormal{dph}})\}$

### v2

@@ 3.1 Increase $\alpha$ to gain power
where the parameter $\alpha^{\textnormal{kn}}\in(0,1]$ can be interpreted as the prespecified initial significance level. The difference of $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}$ to the stopping time $T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}$ of Barber and Candès (2015) (see equation (1)) is the second condition $\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-t\}=0$. Be aware that this condition only comes into play if the first condition is not satisfied for any $t$, meaning $T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}=T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}$ if $R^{\textnormal{BC}}\neq\emptyset$. However, in case of $T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}=\infty$, we have $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}<\infty$. Similar to BC-knockoff, our procedure, which we call post-hoc knockoff (ph-knockoff), rejects all hypotheses $H_{0}^{i}$ with $W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}$. Since $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\leq T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}$, ph-knockoff is uniformly more powerful than BC-knockoff by Barber and Candès (2015). However, in case of $T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}=\infty$, we can only report the rejections at a larger significance level $\alpha>\alpha^{\textnormal{kn}}$, as we show in the next section (see Theorem 2).
Remark 1
Luo et al. (2025) and Ren and Barber (2024) used a modification of the stopping time $T_{\alpha^{\textnormal{kn}}}^{\textnormal{BC}}$ that is similar to $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}$, adding the condition $\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq t\}<1/\alpha^{\textnormal{kn}}$ instead of $\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-t\}=0$. They used this adjusted stopping time to improve their calibrated and derandomized knockoff procedures without considering a post-hoc choice of $\alpha$. While our approach of switching $\alpha$ post-hoc would also work with their stopping time, our stopping time $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}$ usually leads to smaller post-hoc levels and therefore we think $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}$ is more appropriate.
@@ 3.2 Decrease $\alpha$ to improve precision
In Algorithm 1, we summarize our general procedure, where we denote by $\tilde{\alpha}^{\textnormal{ph}}$ our post-hoc significance level and by $R^{\textnormal{ph}}$ the corresponding rejection set. If BC-knockoff does not make any rejections, we increase the significance level ($\tilde{\alpha}^{\textnormal{ph}}>\alpha^{\textnormal{kn}}$), and thus potentially obtain rejections. If BC-knockoff makes rejections, we decrease the significance level ($\tilde{\alpha}^{\textnormal{ph}}\leq\alpha^{\textnormal{kn}}$) to increase the evidence. In this way, the claims made by ph-knockoff are always stronger than the claims made by BC-knockoff.
@@
Output: Rejection set $R^{\textnormal{ph}}$ and post-hoc level $\tilde{\alpha}^{\textnormal{ph}}$.
@@
4:  $\tilde{\alpha}^{\textnormal{ph}}=\frac{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}}{\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}}$
@@
6:  $\tilde{\alpha}^{\textnormal{ph}}=\alpha^{\textnormal{kn}}$
@@
8:return $R^{\textnormal{ph}}$, $\tilde{\alpha}^{\textnormal{ph}}$
@@
The validity of these local e-values (i.e., that $\mathbb{E}_{H_{0}^{S}}[E_{S}^{\textnormal{ph}}]\leq 1$) follows since $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}$ is a stopping time with respect to the filtration generated by a masked version of $(W_{i})_{i\in[p]}$, meaning the event $\{T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\leq t\}$ is determined by the absolute values $|W_{i}|$ for all $i$ and $\sum_{i:|W_{i}|\geq s}\mathbbm{1}\{W_{i}>0\}$ for all $s\leq t$ (see Remark 4 in (Ren and Barber, 2024)).
Theorem 2
@@
$\tilde{\alpha}^{\textnormal{ph}}=\frac{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}}{\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}}\quad\text{with the convention }\tilde{\alpha}^{\textnormal{ph}}=\alpha^{\textnormal{kn}}\text{ if }\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}=0.$
Then $R^{\textnormal{ph}}\in\mathcal{R}_{\tilde{\alpha}^{\textnormal{ph}}}(\boldsymbol{E}^{\textnormal{ph}})$, where $R^{\textnormal{ph}}=\{i\in[p]:W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}$. Therefore, it holds that
$\mathbb{E}\left[\frac{\textnormal{FDP}_{I_{0}}(R^{\textnormal{ph}})}{\tilde{\alpha}^{\textnormal{ph}}}\right]\leq 1.$ (7)
Furthermore, $R^{\textnormal{ph}}\supseteq R^{\textnormal{BC}}$, meaning ph-knockoff always makes as least as many rejections as BC-knockoff and $\tilde{\alpha}^{\textnormal{ph}}\leq\alpha^{\textnormal{kn}}$ on the event $\{R^{\textnormal{BC}}\neq\emptyset\}$, meaning if BC-knockoff rejects any hypothesis then ph-knockoff does not report a larger significance level.
Proof We need to show that $E_{S}^{\textnormal{ph}}\geq\textnormal{FDP}_{S}(R^{\textnormal{ph}})/\tilde{\alpha}^{\textnormal{ph}}$ for all $S\subseteq[p]$. First, note that in case of $\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}=0$, we have $R^{\textnormal{ph}}=\emptyset$ and therefore the condition is trivially fulfilled. Hence, assume $R^{\textnormal{ph}}\neq\emptyset$. Then,
$E_{S}^{\textnormal{ph}}=\frac{|S\cap R^{\textnormal{ph}}|}{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}}=\frac{\textnormal{FDP}_{S}(R^{\textnormal{ph}})|R^{\textnormal{ph}}|}{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}}=\frac{\textnormal{FDP}_{S}(R^{\textnormal{ph}})}{\tilde{\alpha}^{\textnormal{ph}}}.$
Remark 3
Note that $\tilde{\alpha}^{\textnormal{ph}}$ equals 11 in case of $\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}=1$. While rejections at a level of 11 seems like a trivial claim in the fixed-$\alpha$ regime, it is not for random $\alpha$ (Koning, 2023). However, if such large values for $\tilde{\alpha}^{\textnormal{ph}}$ should be avoided, one could just set $\tilde{\alpha}^{\textnormal{ph}}=\alpha^{\textnormal{kn}}$ and $R^{\textnormal{ph}}=\emptyset$ in those cases. This obviously maintains (7).
@@ 4 Post-hoc $\alpha$ with derandomized knockoffs
Ren and Barber (2024) introduced a derandomized knockoff procedure, which we call RB-knockoff in the following, by generating multiple knockoff copies and merging the results using an average of (compound) e-values. In the following, we briefly recap their method and demonstrate how it can be improved by using a post-hoc significance level. It should be noted that we present the RB-knockoff procedure with our stopping time $T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}$ to be consistent with the previous part, even though Ren and Barber (2024) proposed a slightly different one (see Remark 1). However, our approach generally works for any stopping time.
@@
Ren and Barber (2024) showed that $E_{1}^{\text{avg}},\ldots,E_{p}^{\text{avg}}$ are compound e-values, which means $\mathbb{E}\left[\sum_{i\in I_{0}}E_{i}^{\text{avg}}\right]\leq p$. The RB-knockoff procedure is defined by e-BH (Wang and Ramdas, 2022) applied on $E_{1}^{\text{avg}},\ldots,E_{p}^{\text{avg}}$ at prespecified level $\alpha^{\textnormal{ebh}}\in(0,1]$ and controls the FDR at level $\alpha^{\textnormal{ebh}}$.
In case of $k=1$ and $\alpha^{\textnormal{ebh}}=\alpha^{\textnormal{kn}}$, RB-knockoff coincides with BC-knockoff (Ren and Barber, 2024). However, Ren and Barber (2024) showed that $\alpha^{\textnormal{ebh}}=\alpha^{\textnormal{kn}}$ often leads to zero power in case of $k>1$ and therefore recommend to choose $\alpha^{\textnormal{ebh}}>\alpha^{\textnormal{kn}}$. The choice of the parameter $\alpha^{\textnormal{ebh}}$ is crucial. If $\alpha^{\textnormal{ebh}}$ is too small, one might not be able to make any rejections even though the signal is strong. If $\alpha^{\textnormal{ebh}}$ is too large, the results become imprecise. Ren and Barber (2024) derive some concrete choices like $\alpha^{\textnormal{ebh}}=2\alpha^{\textnormal{kn}}$ based on simulations and heuristics, but they do not propose an optimal parameter. We note that it is possible to choose $\alpha^{\textnormal{ebh}}$ completely data-dependent while maintaining (4). In this way, we can just choose the $\alpha^{\textnormal{ebh}}$ that leads to the most desirable rejection set.
Theorem 4
@@
$\mathbb{E}\left[\sup_{\alpha^{\textnormal{ebh}}\in(0,1]}\frac{\textnormal{FDP}_{I_{0}}(R_{\alpha^{\textnormal{ebh}}}^{\textnormal{RB}})}{\alpha^{\textnormal{ebh}}}\right]\leq 1.$
@@
The proof of the Theorem 4 follows immediately from the equivalence of e-BH with compound e-values $E_{1}^{\text{avg}},\ldots,E_{p}^{\text{avg}}$ and the e-Closure Principle with the following local e-values (see Theorem 23 in (Xu et al., 2025))
@@
There are multiple reasonable choices to pick $\alpha^{\textnormal{ebh}}$ based on the data that could lead to different rejection sets (in the following we refer to our data-dependent choice as $\tilde{\alpha}^{\textnormal{dph}}$). In order to compare our derandomized ph-knockoff procedure to RB-knockoff in the following, we start with some initial level $\alpha^{\textnormal{ebh}}\in[0,1]$ that should be chosen as in the RB-knockoff procedure. If RB-knockoff does not make any rejections with that $\alpha^{\textnormal{ebh}}$, we choose $\tilde{\alpha}^{\textnormal{dph}}>\alpha^{\textnormal{ebh}}$ as small as possible such that $R_{\tilde{\alpha}^{\textnormal{dph}}}^{\textnormal{RB}}$ is not empty. If $R_{\alpha^{\textnormal{ebh}}}^{\textnormal{RB}}\neq\emptyset$, then we choose $\tilde{\alpha}^{\textnormal{dph}}\leq\alpha^{\textnormal{ebh}}$ as small as possible while $R_{\tilde{\alpha}^{\textnormal{dph}}}^{\textnormal{RB}}=R_{\alpha^{\textnormal{ebh}}}^{\textnormal{RB}}$. In this way, our derandomized ph-knockoff procedure always rejects as least as many hypotheses as RB-knockoff. Furthermore, if RB-knockoff makes rejections, then $\tilde{\alpha}^{\textnormal{dph}}\leq\alpha^{\textnormal{ebh}}$. Our general procedure is summarized in Algorithm 2.
Remark 5
In case of $k=1$ the derandomized ph-knockoff method (Algorithm 2) and the ph-knockoff procedure (Algorithm 1) are equivalent regardless of the chosen $\alpha_{\textnormal{ebh}}$. To see this, note that for $k=1$ we have
Hence, if $R_{\alpha^{\textnormal{ebh}}}^{\textnormal{RB}}\neq\emptyset$, then $R^{\textnormal{dph}}=R_{\alpha^{\textnormal{ebh}}}^{\textnormal{RB}}=\{i\in[p]:W_{i}\geq T_{\alpha^{\textnormal{kn}}}^{\textnormal{ph}}\}=R^{\textnormal{ph}}$ and $\tilde{\alpha}^{\textnormal{dph}}=p/(i^{\textnormal{ebh}}E_{(i^{\textnormal{ebh}})}^{\text{avg}})=\frac{1+\sum_{i\in[p]}\mathbbm{1}\{W_{i}\leq-T_{{\alpha^{\textnormal{kn}}}}^{\textnormal{ph}}\}}{\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq T_{{\alpha^{\textnormal{kn}}}}^{\textnormal{ph}}\}}=\tilde{\alpha}^{\textnormal{ph}}$. Furthermore, if $R_{\alpha^{\textnormal{ebh}}}^{\textnormal{RB}}=\emptyset$, it holds that $i^{*}=\sum_{i\in[p]}\mathbbm{1}\{W_{i}\geq T_{{\alpha^{\textnormal{kn}}}}^{\textnormal{ph}}\}$, which implies $\tilde{\alpha}^{\textnormal{dph}}=\tilde{\alpha}^{\textnormal{ph}}$ and $R^{\textnormal{dph}}=R^{\textnormal{ph}}$ in the same manner as before. Hence, the post-hoc framework unifies the non-derandomized and the derandomized knockoff method.
@@
Output: Rejection set $R^{\textnormal{dph}}$ and post-hoc level $\tilde{\alpha}^{\textnormal{dph}}$.
@@
5:  $\tilde{\alpha}^{\textnormal{dph}}=p/(i^{\textnormal{ebh}}E_{(i^{\textnormal{ebh}})}^{\text{avg}})$
@@
10:   $\tilde{\alpha}^{\textnormal{dph}}=\alpha^{\textnormal{kn}}$
@@
13:   $\tilde{\alpha}^{\textnormal{dph}}=p/(i^{*}E_{(i^{*})}^{\text{avg}})$
14:   $R^{\textnormal{dph}}=\{i\in[p]:E_{i}^{\text{avg}}\geq p/(i^{*}\tilde{\alpha}^{\textnormal{dph}})\}$